# Extracting self-similarity from data

### v2

@@
Extracting self-similarity from data
Nikos Bempedelis,1,2∗ Luca Magri,2,3,4 Konstantinos Steiros2∗
1School of Engineering and Materials Science,
Queen Mary University of London, E1 4NS, London, UK
2Department of Aeronautics, Imperial College London, SW7 2AZ London, UK
3The Alan Turing Institute, London NW1 2DB, UK
4Politecnico di Torino, DIMEAS, Corso Duca degli Abruzzi, 24 10129 Torino, Italy
Corresponding authors: n.bempedelis@qmul.ac.uk (NB), k.steiros@imperial.ac.uk (KS)
@@
The identification of self-similarity is an indispensable tool for understanding and modelling physical phenomena. Unfortunately, this is not always possible to perform formally in highly complex problems. We propose a methodology to extract the similarity variables of a self-similar physical process directly from data, without prior knowledge of the governing equations or boundary conditions, based on an optimization problem and symbolic regression. We analyze the accuracy and robustness of our method in five problems which have been influential in fluid mechanics research: a laminar boundary layer, Burger’s equation, a turbulent wake, a collapsing cavity, and grid-generated turbulence. Our analysis considers datasets acquired via both numerical and wind-tunnel experiments. The algorithm recovers the known self-similarity expressions in the first four problems and generates new insights in the final problem (grid turbulence).
Introduction
@@
The various constraints that describe a physical phenomenon are often reflections of underlying symmetry principles, summarising regularities that exist independent of specific dynamics  (?). A notable example is the constraint of momentum conservation, which reflects a translational symmetry of the Euler-Lagrange equations  (?). Dimensional analysis, historically linked to the discovery of scaling laws and non-dimensional numbers  (?, ?), expresses the principle of covariance, i.e., the symmetry of any physical law under a dilational transformation of its units of measurement. More generally, when one considers a specific problem, i.e., the governing physical laws and the corresponding boundary and/or initial conditions, multiple symmetries may be simultaneously present (e.g., spiral, reflectional, rotational)  (?, ?).
@@
A type of symmetry which is of fundamental importance across various branches of physics is that of self-similarity  (?, ?, ?, ?). Following Refs.  (?, ?, ?), we refer to self-similar (or self-preserving) phenomena as those whose evolution remains invariant under the group transformations of dilation (e.g., heat diffusion), translation (e.g., travelling waves), or a combination of both (e.g., turbulent wakes). Self-similarity allows for the reduction of an $n-$independent variable partial differential equation system into a system with $n-1$ similarity variables  (?, ?, ?). The transformation from original to similarity variables is known as a similarity transformation. Similarity transformations are particularly useful in problems involving $n=2$ independent variables because they transform a partial differential equation into an ordinary differential equation, which is more amenable to be analytically solved  (?, ?). Even if the governing equations are unknown, similarity considerations allow for the grouping of the problem parameters into similarity variables, reducing the effort required by the experimentalist when performing parametric characterizations of a problem by orders of magnitude  (?). Additionally, the identification of similarity variables, subject to the constraints of the governing equations, can lead to the derivation of scaling laws, which describe the asymptotic evolution of the problem variables  (?, ?, ?). Fluid mechanics is a branch of physics where identification of similarity transformations and variables has played a prominent role. Indeed, self-similarity is at the heart of theoretical efforts for the modelling of laminar and turbulent boundary layers  (?, ?), free shear flows  (?, ?, ?, ?), cascade dynamics  (?, ?, ?, ?), linear and nonlinear waves  (?, ?, ?, ?), singularities  (?), high Mach number aerofoil design  (?), among many others.
@@
There are three possible strategies to identify self-similarity. First, if the differential equations describing the problem are known and the boundary conditions are simple, one can formally extract the flow symmetries using the theory of Lie groups  (?, ?). Second, if the equations are unknown one may still invoke dimensional analysis to uncover dilational symmetries  (?). Finally, on many occasions, human intuition has allowed the uncovering of self-similarity directly, via visual inspection of the problem solution (as for example in the case of turbulent shear flows  (?)). However, this is not always sufficient to reveal the underlying self-similarity of a phenomenon. Intuition depends on the skills of the practitioner and is generally confined to simple problems. Dimensional analysis can only treat problems with dimensional variables, while even then it may uncover only the ‘tip of the iceberg’ of possible self-similar solutions, i.e., some (but far from all) of the self-similarities connected to dilational transformations  (?). Lie group theory may become unfeasible in cases where the boundary conditions are overly complicated. Even if this is not the case, it may reveal some but not all symmetries of the problem. A coordinate transformation may be necessary for additional, ‘hidden’ symmetries to be revealed  (?, ?). In other cases the governing equations may be inadequate. An example of the above occurs in turbulent flows, where it is customary to consider the ensemble-averaged flow equations of motion known as RANS (Reynolds-Averaged Navier-Stokes). Given the chaotic turbulent dynamics and complex boundary conditions, certain original symmetries of the Navier-Stokes equations are broken, but are still hypothesized to be recovered in a statistical sense  (?). However, due to the underdetermined nature of the RANS equations (also known as the turbulence closure problem), extraction of symmetries is challenging  (?).
@@
The recent emergence of machine learning methodologies has provided a significant boost in our ability to uncover symmetries and self-similarities  (?, ?, ?). If self-similarity exists in a problem, it must appear in the observables (data), and can be, in principle, discovered by data-driven methods. For instance, a known challenge in conventional dimensional analysis (Buckingham Pi theorem) is that it yields a non-unique set of non-dimensional variables that govern the evolution of a physical problem  (?, ?). Information from data can in that case provide a constraint to Buckingham Pi, and has been recently used to identify the most appropriate set of non-dimensional variables  (?, ?, ?, ?, ?, ?). Other examples are cases in which conservation laws and symmetries of ordinary differential equations cannot be readily extracted because they are hidden, i.e., they require a coordinate transformation before they become manifest. Machine learning can be used to provide the necessary coordinate transformation for hidden symmetries to manifest  (?, ?). Of high relevance are also efforts to leverage the information contained in data to close underdetermined governing equations  (?), which potentially opens the door for more extended Lie group analysis approaches.
@@
In this work, we present a generalized methodology that can identify if similarity variables exist and, if they do, their mathematical expressions directly from data, without prior knowledge of the governing equations or boundary conditions. Our work differs from previous efforts (e.g.,  (?, ?)) in that it is not based on dimensional analysis (although it is dimensionally consistent) and can thus identify self-similarities beyond dilational transformations connected to dimensional problems. This is achieved by formulating a minimization problem to identify the similarity variables, which are then interpreted analytically via symbolic regression. In particular, our method simultaneously searches for the optimal non-linear transformations of both independent variables (spatiotemporal coordinates) and dependent variables (observables) that yield the similarity variables of the problem and is thus different to methods aiming to identify coordinate transformations that render hidden symmetries manifest (e.g.,  (?)). In the next section and supplementary material we describe the algorithm that extracts the similarity variables of a self-similar physical process from data. The potential utility of the proposed method is demonstrated by applying it to five problems which have been influential in fluid dynamics research and are exactly or approximately self-similar under different transformation types, with data derived from both laboratory and numerical experiments.
Data-driven identification of self-similarity
@@
Let $q(s,t)$ be a quantity of interest governed by a set of non-linear partial differential equations, with $s$ and $t$ being the independent variables, typically associated with a spatial coordinate and time, respectively. In some problems (e.g., a laminar boundary layer), $t$ may also correspond to a spatial coordinate. Consider the set of similarity variables $\xi$ and $\tilde{q}$ which lead to the similarity transformation $q(s,t)\rightarrow\tilde{q}(\xi,t)\equiv\tilde{q}(\xi)$, with $\xi=\xi(s,t)$. $\xi$ may also depend on the governing parameters (constants) of the problem. Assuming that we have measurements of $q(s,t)$ at distinct instants (stations) $t_{i}$, where $i=1,\dots,n_{t}$, we propose a two-stage workflow for extracting the similarity transformation directly from data, without previous knowledge of the governing equations. We note that since the self-similarity can be inexact or the data may include error, the equivalence $\tilde{q}(\xi,t)\equiv\tilde{q}(\xi)$ can be at most expected to hold approximately. In higher dimensions, $s$ and $\xi$ are vectors.
Step 1. Search for similarity variables.
@@
Decomposing the similarity variables in the form of distinct transformations is important for facilitating the success of the ensuing optimization and regression tasks, as well as for enhancing the interpretability of the method. The search for similarity variables is formulated as a minimization problem, {argmini}—s— w 12 ∑_i = 1^n_t ∑_j = 1^n_t ∥ ~q(ξ,t_i) - ~q(ξ,t_j) ∥_2^2 , where $\boldsymbol{w}$ is the design variable matrix containing the discrete values of the transformation functions $\alpha$, $\beta$, $\gamma$, $\delta$. The $l_{2}$-norms, $\|\cdot\|_{2}^{2}$, are computed following interpolation on the transformed coordinates ($\xi$) grid. In some problems, solving Eq. (Step 1. Search for similarity variables.) can lead to degenerate solutions (i.e., $\tilde{q}=0$), in which case Eq. (Step 1. Search for similarity variables.) is replaced by a mean-regularized cost functional, which prevents trivial optima {argmini}—s— w 12 ∑_i = 1^n_t ∑_j = 1^n_t ~q(ξ,ti) - ~q(ξ,tj)— (~q(ξ,ti) + ~q(ξ,tj) )/2— _2^2 .
This step provides the discrete values of the functions $\alpha$, $\beta$, $\gamma$, $\delta$, and by extension through Eq. (1), the discrete values of the similarity variables $\xi$ and $\tilde{q}$, at all $n_{t}$ instants. In practice, in order to further promote the success of the optimization task whilst also preserving the automated character of the developed method, we propose the implementation of step 1 in successive iterations of increased decomposition complexity (i.e., building gradually from simple dilation ($\beta=\delta=0$) to the general decomposition ($\xi=\alpha(t)s+\beta(s,t),\;\tilde{q}=\gamma(t)q+\delta(s,t)$)). Depending on the problem (e.g., its boundary conditions), certain candidate transformations might be irrelevant or inadmissible (for example, translation of radial coordinates in an axisymmetric problem). A pseudocode describing an example implementation of the method can be found in the Supplementary Material.
Step 2. Analytic form of the transformations.
@@
To reduce the complexity of the regression problem, it is beneficial to take advantage of any prior knowledge regarding the characteristics of the dataset. We enforce two properties. The first property that is critical for discovering physical laws is that of dimensional homogeneity: the units of the identified expression $\psi(\boldsymbol{A})$ should match those of the given dataset $\mathcal{D}$. The second property that is relevant to this work is that scale invariance is related to power laws  (?), allowing us in certain cases to restrict the search for expressions in the form of monomials.
[…]
Results
@@
The thin layer of fluid in the immediate vicinity of a solid boundary is known as the boundary layer. Knowledge of the boundary layer properties is of enormous importance for many applications in the transportation and energy sectors. For two-dimensional, laminar and incompressible boundary layers Prandtl showed that some terms of the Navier-Stokes equations are negligible, and reduced the latter to the “Boundary Layer Equations” (BLE)  (?). By utilising the streamfunction $\Psi$, where $\Psi_{y}=u$ and $\Psi_{x}=-v$, with $u$ the streamwise flow velocity and $v$ the crosswise velocity, the BLE can be expressed in a single equation,
$\Psi_{y}\Psi_{xy}-\Psi_{x}\Psi_{yy}-\nu\Psi_{yyy}=0$ (2)
where $x$ and $y$ are the streamwise and normal-to-the-wall distances, respectively, $\nu$ is the kinematic viscosity of the fluid, and the subscripts denote partial differentiation. Blasius  (?) noted that for semi-infinite flat plate boundary conditions the BLE is symmetrical under a dilational transformation, which reduces the problem to an ordinary differential equation for the nondimensional streamfunction $f(\tilde{y})=\Psi/\sqrt{\nu U_{\infty}x}$, with $\tilde{y}=y\sqrt{U_{\infty}/(\nu x)}$. In the above, $U_{\infty}$ is the free stream velocity away from the flat plate. The BLE are thus reduced to the boundary value problem
@@
Knowledge of $\tilde{u}=f_{\tilde{y}}$ (e.g., via numerical integration of the Blasius equation) allows for the calculation of the streamwise velocity distribution $u=U_{\infty}\tilde{u}$ at any location in the boundary layer. Here, we derive the Blasius similarity directly from data, without prior knowledge of the Navier-Stokes equations, Prandtl’s scaling analysis, or Blasius’ similarity arguments. Our dataset consists of $u(x,y)$ data from a flat plate laminar boundary layer flow (Fig. 1(a)), acquired by solving the Blasius ODE with a shooting method  (?). In particular, we extract velocity profiles $u(y)$ from nine stations at different streamwise coordinates $x$ (Fig. 1(b)). The algorithm commences by seeking similarity under the dilational transformations $\tilde{y}=\alpha(x)y$, $\tilde{u}=\beta(x)u$. Figure 1(c) shows that the transformed profiles have collapsed onto a single curve. The identified transformations for the wall-normal coordinate $y$ and the streamwise velocity $u$ are shown in figures 1(d) and (e), respectively.
The second step consists of expressing the identified transformations as functions of the governing parameters and independent variables, i.e., a symbolic regression task. The library of candidate variables is composed of the free-stream velocity $U_{\infty}$, the viscosity $\nu$, and the streamwise coordinate of the extracted profiles $x$, i.e., $\alpha=\alpha(U_{\infty},\nu,x)$ and $\beta=\beta(U_{\infty},\nu,x)$. Owing to the nature of the assumed dilational transformations, we look for expressions in the form of monomials. Dimensional homogeneity is enforced by assuming $\tilde{y}$ and $\tilde{u}$ to be dimensionless and a non-zero weighting constant $w_{D}$ in the objective function (Eq. (Step 2. Analytic form of the transformations.)). The interpreted expressions are
@@
which are in close agreement with the theoretically derived similarity variables of Blasius ($\tilde{y}=y\sqrt{U_{\infty}/(\nu x)}$, $\tilde{u}=u/U_{\infty}$). We note that the identified scalings can vary by an arbitrary multiplicative constant (Fig. 1(d)), depending on whether and how the user normalises the collapsed profiles.
@@
Burgers’ equation is a partial differential equation that finds wide application in fluid dynamics, non-linear acoustics and traffic flow among others. It acts as a prototype for a variety of phenomena including shock wave formation, rarefaction waves, and turbulence  (?, ?). In the absence of diffusion, it is the simplest model for gasdynamics. Its mathematical expression in one dimension, along with its initial condition is

### v3

@@
Abstract
@@
Identifying self-similarity is key to understanding and modelling a plethora of phenomena in fluid mechanics. Unfortunately, this is not always possible to perform formally in highly complex flows. We propose a methodology to extract the similarity variables of a self-similar physical process directly from data, without prior knowledge of the governing equations or boundary conditions, based on an optimization problem and symbolic regression. We analyze the accuracy and robustness of our method in five problems which have been influential in fluid mechanics research: a laminar boundary layer, Burger’s equation, a turbulent wake, a collapsing cavity, and decaying turbulence. Our analysis considers datasets acquired via both numerical and wind tunnel experiments. The algorithm recovers the known self-similarity expressions in the first four problems and generates new insights on single length scale theories of homogeneous turbulence.
keywords:
1 Introduction
@@
The various constraints that describe a physical phenomenon are often reflections of underlying symmetry principles, summarising regularities that exist independent of specific dynamics (Gross, 1996). A notable example is the constraint of momentum conservation, which reflects a translational symmetry of the Euler-Lagrange equations (Landau & Lifshitz, 1976). Dimensional analysis, historically linked to the discovery of scaling laws and non-dimensional numbers (Barenblatt, 1996; Cantwell, 2002), expresses the principle of covariance, i.e., the symmetry of any physical law under a dilational transformation of its units of measurement. More generally, when one considers a specific problem, i.e., the governing physical laws and the corresponding boundary and/or initial conditions, multiple symmetries may be simultaneously present (e.g., spiral, reflectional, rotational) (Pakdemirli & Yurusoy, 1998; Cantwell, 2002).
@@
A type of symmetry which is of fundamental importance across various branches of physics is that of self-similarity (Barenblatt, 1996; Cantwell, 2002; Tennekes & Lumley, 1972; Townsend, 1976). Following Barenblatt (1996); Townsend (1976); Pope (2000), we refer to self-similar (or self-preserving) phenomena as those whose evolution remains invariant under the group transformations of dilation (e.g., heat diffusion), translation (e.g., travelling waves), or a combination of both (e.g., turbulent wakes). Self-similarity allows for the reduction of an $n$-independent variable partial differential equation system into a system with n1n-1 similarity variables (Pakdemirli & Yurusoy, 1998; Cantwell, 2002; Birkhoff, 1960). The transformation from original to similarity variables is known as a similarity transformation. Similarity transformations are particularly useful in problems involving $n=2$ independent variables because they transform a partial differential equation into an ordinary differential equation, which is more amenable to be analytically solved (Pakdemirli & Yurusoy, 1998; Cantwell, 2002). Even if the governing equations are unknown, similarity considerations allow for the grouping of the problem parameters into similarity variables, reducing the effort required by the experimentalist when performing parametric characterizations of a problem by orders of magnitude (Barenblatt, 1996). Additionally, the identification of similarity variables, subject to the constraints of the governing equations, can lead to the derivation of scaling laws, which describe the asymptotic evolution of the problem variables (Tennekes & Lumley, 1972; Townsend, 1976; Beaumard et al., 2024). The identification of similarity transformations and variables has played a prominent role in fluid mechanics research. Indeed, self-similarity is at the heart of theoretical efforts for the modelling of laminar and turbulent boundary layers (Prandtl, 1904; Townsend, 1976), free shear flows (Tennekes & Lumley, 1972; Townsend, 1976; Pope, 2000; George, 1989), cascade dynamics (Kolmogorov, 1941b; Pope, 2000; Vassilicos, 2015; Steiros, 2022a), linear and nonlinear waves (Taylor, 1950a, b; Zel’Dovich & Raizer, 1967; Whitham, 2011), singularities (Eggers & Fontelos, 2015), high Mach number aerofoil design (Cantwell, 2002), among many others.
@@
There are three possible strategies to identify self-similarity. First, if the differential equations describing the problem are known and the boundary conditions are simple, one can formally extract the flow symmetries using the theory of Lie groups (Cantwell, 2002; Birkhoff, 1960; Cantwell, 1978). Second, if the equations are unknown one may still invoke dimensional analysis to uncover dilational symmetries (Barenblatt, 1996). Finally, on many occasions, intuition has allowed the uncovering of self-similarity directly, via visual inspection of the problem solution (as for example in the case of turbulent shear flows (Townsend, 1976)). However, this is not always sufficient to reveal the underlying self-similarity of a phenomenon. Intuition depends on the skills of the practitioner and is generally confined to simple problems. Dimensional analysis can only treat problems with dimensional variables, while even then it may uncover only the ‘tip of the iceberg’ of possible self-similar solutions, i.e., some (but far from all) of the self-similarities connected to dilational transformations (Barenblatt, 1996). Lie group theory may become unfeasible in cases where the boundary conditions are overly complicated. Even if this is not the case, it may reveal some but not all symmetries of the problem. A coordinate transformation may be necessary for additional, ‘hidden’ symmetries to be revealed (Cantwell, 2002; Liu & Tegmark, 2022). In other cases the governing equations may be inadequate. An example of the above occurs in turbulent flows, where it is customary to consider the ensemble-averaged flow equations of motion (i.e. the Reynolds-Averaged Navier-Stokes equations). Given the chaotic turbulent dynamics and complex boundary conditions, certain original symmetries of the Navier-Stokes equations are broken, but are still hypothesized to be recovered in a statistical sense (Frisch, 1995). However, due to the underdetermined nature of the RANS equations (turbulence closure problem), extraction of symmetries is challenging (Oberlack, 2001, 1999; Oberlack et al., 2006; Oberlack & Rosteck, 2010; Oberlack et al., 2022).
@@
The recent emergence of machine learning methodologies has provided a significant boost in our ability to uncover symmetries and self-similarities (Desai et al., 2022; Yang et al., 2023; Otto et al., 2023). If self-similarity exists in a problem, it must appear in the observables (data), and can be, in principle, discovered by data-driven methods. For instance, a known challenge in conventional dimensional analysis (Buckingham Pi theorem) is that it yields a non-unique set of non-dimensional variables that govern the evolution of a physical problem (Barenblatt, 1996; Cantwell, 2002). Information from data can in that case provide a constraint to Buckingham Pi, and has been recently used to identify the most appropriate set of non-dimensional variables (Mendez & Ordóñez, 2004; Constantine et al., 2017; Jofre et al., 2020; Saha et al., 2021; Xie et al., 2022; Bakarji et al., 2022; Yuan & Lozano-Durán, 2025). Other examples are cases in which conservation laws and symmetries of ordinary differential equations cannot be readily extracted because they are hidden, i.e., they require a coordinate transformation before they become manifest. Machine learning can be used to provide the necessary coordinate transformation for hidden symmetries to manifest (Liu & Tegmark, 2022; Mototake, 2023). Of high relevance are also efforts to leverage the information contained in data to reveal scale-invariant flow structures (Fukami et al., 2024) or close underdetermined governing equations (Duraisamy et al., 2019), which can support the training and testing of nonlinear machine-learning techniques (e.g., Fukami & Taira (2024); Duraisamy et al. (2025)).
@@
In this work, we present a generalized methodology that can identify if similarity variables exist and, if they do, their mathematical expressions directly from data, without prior knowledge of the governing equations or boundary conditions. Our work differs from previous efforts (e.g., Xie et al. (2022); Bakarji et al. (2022); Yuan & Lozano-Durán (2025)) in that it is not based on dimensional analysis (although it is dimensionally consistent) and can thus identify self-similarities beyond dilational transformations connected to dimensional problems. This is achieved by formulating a minimization problem to identify the similarity variables, which are then interpreted analytically via symbolic regression. In particular, our method simultaneously searches for the optimal non-linear transformations of both independent variables (spatiotemporal coordinates) and dependent variables (observables) that yield the similarity variables of the problem and is thus different to methods aiming to identify coordinate transformations that render hidden symmetries manifest (e.g., Liu & Tegmark (2022)). The paper is structured as follows: in section 2 and Appendix A we describe the algorithm that extracts the similarity variables of a self-similar physical process from data. Section 3 demonstrates the potential utility of the proposed method by applying it to five problems which have been influential in fluid dynamics research and are exactly or approximately self-similar under different transformation types, with data derived from both laboratory and numerical experiments. Finally, Section 4 discusses potential applications and limitations of the method and provides concluding remarks.
2 Data-driven identification of self-similarity
@@
Let $q(s,t)$ be a quantity of interest governed by a set of non-linear partial differential equations, with $s$ and $t$ being the independent variables, typically associated with a spatial coordinate and time, respectively. In some problems (e.g., a laminar boundary layer), $t$ may also correspond to a spatial coordinate. Consider the set of variables $\xi$ and $\tilde{q}$, which lead to the transformation $q(s,t)\rightarrow\tilde{q}(\xi,t)$, with $\xi=\xi(s,t)$, where $\xi$ may also depend on the parameters (constants) of the problem. If $\tilde{q}(\xi,t)$ is independent of $t$, i.e., there is a function $\hat{q}(\xi)$ such that $\tilde{q}(\xi,t)\equiv\hat{q}(\xi)$, then $q(s,t)$ is said to be self-similar (Pope, 2000). In this case, the variables and the transformation are referred to as similarity variables and similarity transformation, respectively. Assuming that we have measurements of $q(s,t)$ at distinct instants (stations) $t_{i}$, where $i=1,\dots,n_{t}$ and $n_{t}$ is the number of available instants (stations), we propose a two-stage workflow for extracting the similarity transformation directly from data, without previous knowledge of the governing equations. Because self-similarity can be inexact or the data may include errors, the equivalence $\tilde{q}(\xi,t)\equiv\hat{q}(\xi)$ can be at most expected to hold approximately. In higher dimensions, $s$ and $\xi$ are vectors.
Step 1. Search for similarity variables
@@
Decomposing the similarity variables in the form of distinct transformations is important for facilitating the success of the ensuing optimization and regression tasks, as well as for enhancing the interpretability of the method. The search for similarity variables is formulated as a minimization problem,
$\underset{\displaystyle\boldsymbol{w}}{\mathrm{arg~min}}\quad\frac{1}{2}\sum_{i=1}^{n_{t}}\sum_{j=1}^{n_{t}}\|\tilde{q}(\xi,t_{i})-\tilde{q}(\xi,t_{j})\|_{2}^{2}\hfil\hfil\hfil\hfil,$ (2)
where $\boldsymbol{w}$ is the design variable matrix containing the discrete values of the transformation functions $\alpha$, $\beta$, $\gamma$, $\delta$. The $l_{2}$-norms, $\|\cdot\|_{2}^{2}$, are computed following interpolation on the transformed coordinates ($\xi$) grid. The resolution of the $\xi$ grid, i.e., the number of points where the $l_{2}$ norms are evaluated, can be freely chosen by the user and can be set equal to the number of points in the input (non-transformed) data (see also Appendix B). In some problems, solving Eq. (2) can lead to degenerate solutions (i.e., $\tilde{q}=0$), in which case Eq. (2) is replaced by a mean-regularized cost functional, which prevents trivial optima
$\underset{\displaystyle\boldsymbol{w}}{\mathrm{arg~min}}\quad\frac{1}{2}\sum_{i=1}^{n_{t}}\sum_{j=1}^{n_{t}}\bigg\|\frac{\tilde{q}(\xi,t_{i})-\tilde{q}(\xi,t_{j})}{\sqrt{|\left(\tilde{q}(\xi,t_{i})+\tilde{q}(\xi,t_{j})\right)/2|}}\bigg\|_{2}^{2}\hfil\hfil\hfil\hfil.$ (3)
This step provides the discrete values of the functions $\alpha$, $\beta$, $\gamma$, $\delta$, and by extension through Eq. (1), the discrete values of the similarity variables $\xi$ and $\tilde{q}$, at all $n_{t}$ instants. In practice, in order to further promote the success of the optimization task whilst also preserving the automated character of the developed method, we propose the implementation of step 1 in successive iterations of increased decomposition complexity (i.e., building gradually from simple dilation ($\beta=\delta=0$) to the general decomposition ($\xi=\alpha(t)s+\beta(s,t),\;\tilde{q}=\gamma(t)q+\delta(s,t)$)). Depending on the problem (e.g., its boundary conditions), certain candidate transformations might be irrelevant or inadmissible (for example, translation of radial coordinates in an axisymmetric problem). Appendix A provides the pseudocode describing an example implementation of the method with clarifying remarks.
Step 2. Analytic form of the transformations
@@
To reduce the complexity of the regression problem, it is beneficial to take advantage of any prior knowledge regarding the characteristics of the dataset. We enforce two properties. The first property that is critical for discovering physical laws is that of dimensional homogeneity: the units of the identified expression $\psi(\boldsymbol{A})$ should match those of the given dataset $\mathcal{D}$. The second property that is relevant to this work is that scale invariance is related to power laws (Barenblatt, 1996), allowing us in certain cases to restrict the search for expressions in the form of monomials.
This step of the algorithm can be carried out with any symbolic regression method, giving users the flexibility to choose their preferred tool. Here, we demonstrate the workflow with two different symbolic regression methods: the open-source library PySR (https://astroautomata.com/PySR/), which offers high performance, flexibility, configurability and generality (Cranmer, 2023), and a custom, simple regression algorithm for the identification of monomials. In both cases, we use an $l_{2}$ loss function, with an optional penalization term enforcing dimensional homogeneity
$\underset{\displaystyle\psi}{\mathrm{arg~min}}\quad\|\psi(\boldsymbol{A})-\mathcal{D}\|_{2}^{2}+w_{D}\|\left[\psi(\boldsymbol{A})\right]-\left[\mathcal{D}\right]\|\hfil\hfil\hfil\hfil,$ (4)
where $w_{D}$ is a non-negative regularization factor and the brackets $\left[\;\right]$ denote the units of a quantity in the form of a dimension vector (see Appendix A for an example). Equation (4) is individually applied to each transformation function ($\alpha$, $\beta$, $\gamma$, or $\delta$). This step provides the analytic form of the similarity transformations $\alpha$, $\beta$, $\gamma$, $\delta$, and thereby of the similarity variables $\xi$ and $\tilde{q}$ through Eq. (1).
3 Results
@@
The first validation example considers the well-known case of the two-dimensional, laminar and incompressible boundary layer, where Prandtl (1904) showed that some terms of the Navier-Stokes equations are negligible, and reduced the latter to the “Boundary Layer Equations” (BLE). By utilising the streamfunction $\Psi$, the BLE can be expressed in a single equation. Blasius (1908) noted that for semi-infinite flat plate boundary conditions the BLE is symmetrical under a dilational transformation, which reduces the problem to an ordinary differential equation for the nondimensional streamfunction $f(\tilde{y})=\Psi/\sqrt{\nu U_{\infty}x}$, with $\tilde{y}=y\sqrt{U_{\infty}/(\nu x)}$. In the above, $U_{\infty}$ is the free stream velocity away from the flat plate, $x$ and $y$ are the streamwise and normal-to-the-wall distances, respectively, and $\nu$ is the kinematic viscosity of the fluid. The BLE are thus reduced to the boundary value problem
@@
Knowledge of $\tilde{u}=f_{\tilde{y}}$ (e.g., via numerical integration of the Blasius equation) allows for the calculation of the streamwise velocity distribution $u=U_{\infty}\tilde{u}$ at any location in the boundary layer. Here, we derive the Blasius similarity directly from data, without prior knowledge of the Navier-Stokes equations, Prandtl’s scaling analysis, or Blasius’ similarity arguments. Our dataset consists of $u(x,y)$ data from a flat plate laminar boundary layer flow (figure 1(a)), acquired by solving the Blasius ODE with a shooting method (Bakarji et al., 2022). In particular, we extract velocity profiles $u(y)$ from nine stations at different streamwise coordinates $x$ (figure 1(b)). The algorithm commences by seeking similarity under the dilational transformations $\tilde{y}=\alpha(x)y$, $\tilde{u}=\beta(x)u$. Figure 1(c) shows that the transformed profiles have collapsed onto a single curve. The identified transformations for the wall-normal coordinate $y$ and the streamwise velocity $u$ are shown in figures 1(d) and (e), respectively.
The second step consists of expressing the identified transformations as functions of the governing parameters and independent variables, i.e., a symbolic regression task. The library of candidate variables is composed of the free-stream velocity $U_{\infty}$, the viscosity $\nu$, and the streamwise coordinate of the extracted profiles $x$, i.e., $\alpha=\alpha(U_{\infty},\nu,x)$ and $\beta=\beta(U_{\infty},\nu,x)$. Owing to the nature of the assumed dilational transformations, we look for expressions in the form of monomials, i.e., $\alpha=c_{1}U_{\infty}^{c_{2}}\nu^{c_{3}}x^{c_{4}}$ and $\beta=c_{5}U_{\infty}^{c_{6}}\nu^{c_{7}}x^{c_{8}}$, where $c_{1}$ and $c_{5}$ are multiplicative constants that reflect the arbitrariness of the dilation factors between the original and similarity variables. Dimensional homogeneity is enforced by assuming $\tilde{y}$ and $\tilde{u}$ to be dimensionless and a non-zero weighting constant $w_{D}$ in the objective function (Eq. (4)). The interpreted expressions are
@@
which are in close agreement with the theoretically derived similarity variables of Blasius ($\tilde{y}=y\sqrt{U_{\infty}/(\nu x)}$, $\tilde{u}=u/U_{\infty}$). As previously noted, the identified and theoretical scalings can vary by an arbitrary multiplicative constant (i.e., the offset in figure 1(d)), depending as to whether and how the user normalizes the collapsed profiles. Appendix B shows results from a sensitivity and robustness analysis with respect to the input data, parameters, and noise.
@@
The second validation example considers Burgers’ equation, which is a partial differential equation that finds wide application in fluid dynamics, non-linear acoustics, and traffic flow, among others. It acts as a prototype for a variety of phenomena including shock wave formation, rarefaction waves, and turbulence (Zel’Dovich & Raizer, 1967; Whitham, 2011). In the absence of diffusion, it is the simplest model for gasdynamics. Its mathematical expression in one dimension, along with its initial condition is